OpenStudy (anonymous):

Find the common ratio for the following sequence. 1/2, -1/4, 1/8, -1/16, ...

OpenStudy (anonymous):

i guess lol i dont know what that means

OpenStudy (anonymous):

the lesson title is geometric sequences

OpenStudy (anonymous):

a sequence in which the next term is found by multiplying the previous term by a constant (the common ratio r)

OpenStudy (luigi0210):

The common ratio is -1/2, since each term is being multiplied by a factor of -1/2

OpenStudy (luigi0210):

We're just looking for the ratio Jhan :P

OpenStudy (anonymous):

IM GETTING CONFUSED lol is it -1/2 or just 1/2

OpenStudy (luigi0210):

Hm, you sure? it does have a pattern of +, -, +, -,...

OpenStudy (anonymous):

both of your answers are possible choices and because i have no idea what im doing it makes it really hard to pick who is right lol

OpenStudy (jhannybean):

Oh, then if we were to avoid the equation then it would just be -1/2 as the common ratio... let's take it two numbers at a time. \[\frac{-1/4}{1/2}=-\frac{1}{2}\]\[\frac{1/8}{-1/4}= -\frac{1}{2}\]\[\frac{-1/16}{1/8}=\frac{-1}{2}\]

OpenStudy (jhannybean):

So yes... the ratio is -1/2...

OpenStudy (luigi0210):

She's great at explanations ^_^

OpenStudy (jhannybean):

For future reference, if you divide the preceding term with the one before, and find a common number between all the divisions, then the common ratio is result from dividing preceding/number before. This applies to geometric sequences.

OpenStudy (anonymous):

i sorta understand now, thank you both again for all the help

OpenStudy (luigi0210):

I wasn't much help.. I just gave an answer D:

OpenStudy (luigi0210):

but you are welcome :P

OpenStudy (anonymous):

lol